# FLORES: A Reconfigured Wheel-Legged Robot for Enhanced Steering and Adaptability

### v1

@@ III-B3 Sim2Real Transfer for FLORES
Before deploying the RL policy on FLORES, we contend with performance degradation caused by discrepancies in physics engines, dynamic modeling, and environmental configurations. To mitigate this sim-to-real gap, we employ domain randomization in Isaac Gym [30], where key parameters such as joint friction, mass distribution, and dynamic properties are randomized (see Table II), ensuring robust performance under diverse conditions. Additionally, we validate and fine-tune the policy in MuJoCo [31], thereby preserving the policy’s superior performance in a new simulation environment while reducing potential risks in real-world experiments.
We further incorporate extended domain randomization to address uncertainties arising from actuators and sensors. For instance, random perturbations are introduced to the motor strength and the PD controller to accommodate varying hardware deviations. We fix the communication frequency between the motor drive and PC at 4000 Hz to maintain precise synchronization of motor commands and smooth actions, and randomize external disturbance forces and push velocities to emulate potential real-world interactions. Moreover, variable observation delays are added to reflect latencies commonly present in sensor processing, thus significantly enhancing the policy’s robustness in real-world scenarios. Owing to these multi-faceted domain randomization strategies, our wheel-legged robot achieves a seamless transition from simulation to reality, maintaining robust performance even in complex and unstructured environments.
TABLE II: Domain Randomization Parameters
Parameter Range Operator
Payload Mass [2, 6](kg) Additive
COM Displacement [-0.2, 0.2](m) Additive
Friction Coefficient [0.6, 2.0] Additive
Motor Strength [0.8, 1.2] Scaling
Proportional Gain ($k_{p}$) [0.9, 1.1] Scaling
Derivative Gain ($k_{d}$) [0.9, 1.1] Scaling
Initial Joint Position [0.8, 1.2] Scaling
Disturbance [-30.0, 30.0](N) Additive
Push Velocity ($x$ $y$) [-1.0,1.0](m/s) Additive
Observation Delay [0,4](steps) Additive
@@ IV Experiment
In this section, we present a comprehensive series of experiments to demonstrate the enhanced steering capabilities, improved navigation efficiency, and versatile locomotion across various terrains. We first show FLORES has the ability to navigate through various terrains with multi-modal locomotion strategies that facilitate smooth transitions between wheeled and legged movement. Additionally, we compute the Cost of Transport (CoT) for FLORES and B2W, a state-of-the-art wheel-legged platform from Unitree Robotics, during various navigation tasks to quantify FLORES’s superior steering ability and navigation efficiency.
@@ IV-A Performance on Various Terrains
Figure 4: The top figure illustrates Flores climbing up the stairs, while the bottom figure depicts the joint angles of the thigh and calf joints for both the front-left and rear-left legs.
FLORES demonstrates its capability for efficient locomotion across diverse terrains in real world, as illustrated in Fig. 4 and Fig. 5. The evaluation encompasses unstructured terrains, including discrete ground (Fig. 5(a)), gravel (Fig. 5(b)), paved surfaces (Fig. 5(c)), grassland (Fig. 5(d)), and slopes (Fig. 5(f)), as well as stairs with a maximum height of 15 cm (Fig. 5(e)). Fig. 4 provides a visual representation of the robot’s joint angle paths for the front thigh and calf joints, illustrating its response when encountering stairs.
@@
Figure 5: By utilizing the modified hip-yaw degree of freedom, FLORES demonstrates efficient and versatile locomotion across various terrains, including (a) discrete terrain, (b) gravel, (c) paved surfaces, (d) grassland, (e) stairs, and (f) slopes. Supplementary videos are available in our project website.
@@ IV-B Comparison of CoT
The primary objective of our experiments is to validate the enhanced steering ability and navigation efficiency of FLORES across various scenarios, including straight-line motion, turning maneuvers, and complex tasks requiring continuous directional adjustments. These scenarios are selected to reflect common operational conditions for mobile robots in unstructured terrains.
Additionally, the data collected from B2W will serve as a benchmark for comparison. By contrasting the performance metrics of FLORES with those of B2W, we aim to highlight the advantages of our robot in terms of energy efficiency and maneuverability. This comparative analysis will underscore the effectiveness of FLORES in navigating diverse environments, further demonstrating its capabilities in real-world applications.
@@
where $\tau$ and $\dot{\theta}$ are the torque and speed of the joint, $[\cdot]^{+}$ is the positive term of the product, $mg$ is the weight of the robot, and $|v^{b}_{xy}|$ is the horizontal speed of the robot’s base. In indoor experiments, $|v^{b}_{xy}|$ is obtained from the motion capture system provided by OptiTrack, while in outdoor environments, it is calculated by averaging the speed during traversal.
@@ IV-B1 Straight-Line Motion
We conduct the straight-line motion experiments on the paved ground, grassland, discrete ground, and gravel terrain with FLORES and B2W. The robots are evaluated at speeds ranging from [0.5 m/s, 1.5 m/s]. The joint torque $\tau$ and joint velocity $\dot{\theta}$ are directly got from the joint encoder. When navigating unstructured terrain, the presence of the modified hip-yaw DoF enables FLORES to achieve this terrain adaptability with minimal energy expenditure, providing almost zero torque through small displacement adjustments during straight-line locomotion. In contrast, the typical hip-roll joint needs to continuously provide torque to maintain the leg at the proper position. The resulting CoT is shown in Fig. 6, where FLORES achieves the lowest CoT on all unstructured terrains.
@@ IV-B2 Lateral Walking
[…]
@@ IV-B3 Turning Maneuvers
[…]
@@ IV-B4 Path Following
Following the validation of simple straight-line and turning motions, we proceed to test a complex path that combines both types of motion to evaluate the navigation efficiency of FLORES. The robot is required with maneuvering through six cluttered boxes, as depicted in Fig. 9. At points A, B, C, D, E, F, and G, the robot is required to make 90 degrees turns. Fig. 9(b) and (c) illustrate the instantaneous CoT for both FLORES and B2W while following the path. Notably, at each turning point, FLORES’s CoT was approximately 70% of that of B2W, and this highlights the superior efficiency of FLORES in navigating through the cluttered environment.
@@
TABLE III: Comparison of CoT for turning maneuvers
@@ V CONCLUSIONS
This letter presents a novel wheel-legged robot, FLORES, featuring an innovative front-leg design that synthesizes the speed and efficiency of wheels with the versatility and adaptability of legs, which significantly enhances the performance on unstructured terrain. FLORES offers enhanced flexibility, improved impact resistance, and faster actuator speeds, making it well-suited for dynamic hybrid locomotion in diverse scenarios. Additionally, we propose an RL-based control framework, with a customized reward mechanism tailored to the unique mechanical design of FLORES. Through real-world experiments, FLORES demonstrates noteworthy flexibility and efficiency across various environments, validating the effectiveness of the proposed design and control strategy.

### v2

@@ III-B3 Sim2Real Transfer for FLORES
Before deployment, we address performance degradation caused by mismatches in physics engines, dynamic models, and environment configurations. To mitigate this sim-to-real gap, we apply domain randomization in Isaac Gym [22] to promote robustness under diverse conditions. We then validate and fine-tune the policy in MuJoCo [30], preserving its performance in a new simulation environment while reducing risks in real-world experiments.
We further extend domain randomization to actuator and sensor uncertainties: we perturb motor strength and PD gains to accommodate hardware deviations, randomize external pushes (forces/velocities) to emulate interactions, and inject variable observation delays to reflect perception latency. The communication frequency between the motor driver and PC is fixed at 4 kHz to stabilize command timing and smooth actions. Together, these strategies narrow the sim-to-real gap for our wheel-legged robot and maintain robust performance even in complex, unstructured environments.
@@ IV Experiment
In this section, we present a comprehensive series of experiments to demonstrate the enhanced steering capabilities, improved navigation efficiency, and versatile locomotion across various terrains. We first show FLORES has the ability to navigate through various terrains with multi-modal locomotion with smooth transitions between wheeled and legged movement. Additionally, we compute the Cost of Transport (CoT) for FLORES and B2W, a state-of-the-art wheel-legged platform from Unitree Robotics, during various navigation tasks to quantify FLORES’s superior steering ability and navigation efficiency. All comparisons use identical task definitions and command profiles.
@@ IV-A Performance on Various Terrains
Figure 4: The top figure illustrates the progress of FLORES as it climbs up the stairs, while the bottom figure depicts the joint angles of the thigh and calf joints for both the front-left and rear-left legs.
FLORES demonstrates efficient real-world locomotion across diverse terrains, as illustrated in Fig. 4 and Fig. 5. The evaluation encompasses unstructured terrains, including discrete ground (Fig. 5(a)), gravel (Fig. 5(b)), paved surfaces (Fig. 5(c)), grassland (Fig. 5(d)), and slopes (Fig. 5(f)), as well as stairs with a maximum height of 15 cm (Fig. 5(e)). Fig. 4 shows the robot’s joint angle paths for the front thigh and calf joints, illustrating its response when encountering stairs.
@@
Figure 5: By utilizing the modified hip-yaw DOF, FLORES demonstrates efficient and versatile locomotion across various terrains, including (a) discrete terrain, (b) gravel, (c) paved surfaces, (d) grassland, (e) stairs, and (f) slopes. Supplementary videos are available on our project website.
@@ IV-B Comparison of CoT
Our experiments validate the enhanced steering ability and navigation efficiency of FLORES across various scenarios, including straight-line motion, turning maneuvers, and complex tasks requiring continuous directional adjustments. These scenarios reflect common operating conditions for mobile robots in unstructured terrains.
Additionally, the data collected from B2W serves as the benchmark. By contrasting the performance metrics of FLORES with those of B2W, we aim to highlight the advantages of our robot in terms of energy efficiency and maneuverability. This comparison underscores the effectiveness of FLORES in navigating diverse environments, further demonstrating its capabilities in real-world applications.
@@
where $\tau$ and $\dot{\theta}$ are the torque and speed of the joint, $\cdot$$^{+}$ is the positive term of the product, $mg$ is the weight of the robot, and $|v^{b}_{xy}|$ is the horizontal speed of the robot’s base. In indoor experiments, $|v^{b}_{xy}|$ is obtained from the motion capture system provided by OptiTrack, while in outdoor environments, it is calculated by averaging the speed during traversal. The B2W wheel modules exhibit an additional Coulomb-like friction torque offset $\tau_{\mathrm{seal}}$ induced by the waterproof sealing structure. To avoid biasing the wheel-side power, we subtract this sealing-induced offset before computing CoT. Let $\mathcal{W}$ denote the set of wheel actuators, and the corrected CoT is defined as
$\mathrm{CoT}_{\text{corr}}=\frac{\sum_{j\notin\mathcal{W}}[\tau_{j}\dot{\theta}_{j}]^{+}\;+\;\sum_{j\in\mathcal{W}}\bigl[\tau_{j}\dot{\theta}_{j}-\tau_{\mathrm{seal}}\lvert\dot{\theta}_{j}\rvert\bigr]^{+}}{mg\,\lvert v^{b}_{xy}\rvert}$. (3)
Note that $\tau_{\mathrm{seal}}$ is measured quasi-statically with a calibrated torque wrench. Unless otherwise specified, all B2W CoT values reported below are computed with the above wheel-sealing friction compensation applied to the wheel joints, while FLORES and all non-wheel joints use the uncorrected term in (2).
@@ IV-B1 Straight-Line Motion
We conduct straight-line motion experiments on the paved ground, grassland, discrete ground, and gravel terrain with FLORES and B2W. The robots are evaluated at speeds ranging from [0.5 m/s, 1.5 m/s]. The joint torque $\tau$ and joint velocity $\dot{\theta}$ are directly obtained from the joint encoder. During planar straight-line tracking, where the robot is required to maintain a consistent heading while rolling forward, the front hip-yaw joint stays almost stationary and applies small corrective torques as needed; the magnitude and frequency of these pulses increase with terrain disturbance. In contrast, the typical hip-roll joint needs to continuously provide torque to maintain the leg at the proper position. The resulting CoT is shown in Fig. 6, where FLORES achieves the lowest CoT on all unstructured terrains.
@@ IV-B2 Lateral Walking
Comparisons of locomotion and CoT between FLORES and B2W for lateral walking are illustrated in Fig. 7. The modification enables the leg structure to rotate so that the front wheel can directly move in the lateral direction, while the rear legs continue to function in a typical lifting motion to facilitate movement. In contrast, the B2W robot requires all legs to be lifted and lowered continuously during lateral movement, which negatively affects the efficiency of its wheeled structure. We test lateral movement for both FLORES and B2W at a commanded speed of $0.5\,\text{m/s}$. FLORES achieves a CoT of 0.3614, which is significantly superior to B2W’s CoT of 0.6767.
@@ IV-B3 Turning Maneuvers
To validate the effectiveness of FLORES’s steering capabilities, we compare the CoT between FLORES and B2W while following circular paths of varying radius. As shown in Fig. 8, the robots follow circular paths with a radius ranging from [0.5 m, 2 m], and travel at a constant input forward velocity $v_{x}=0.4\,\textup{m/s}$. Using measured yaw from OptiTrack, the PID controller maps the tracking error to the desired angular velocity $\omega_{z}^{\text{cmd}}$ for the RL policy. During these path-following tasks, FLORES utilizes the modified hip-yaw DoF as its steering mechanism to adjust its heading direction. In this maneuver, only the rear legs of FLORES need to be lifted while the front legs remain in contact with the ground. In contrast, B2W needs to lift all its legs to continuously change its heading direction, resulting in significantly higher energy expenditure. Table II illustrates that the smaller the turning radius, the more pronounced the advantages of FLORES become in terms of efficiency. This is because a smaller radius necessitates more rapid changes in the robot’s heading direction. FLORES can accommodate this requirement effectively by simply using a larger steering angle provided by its modified hip-yaw DoF. However, B2W must lift and lower its legs at a higher frequency, which leads to increased energy consumption and reduced efficiency during turns.
@@ IV-B4 Path Following
After validating straight-line and turning motions, we test a complex path combining both motions to evaluate the navigation efficiency of FLORES. The robot is required with maneuvering through six cluttered boxes, as depicted in Fig. 9. At points A, B, C, D, E, F, and G, the robot is required to make 90 degrees turns. Fig. 9(b) and (c) illustrate the instantaneous CoT for both FLORES and B2W while following the path. Notably, at each turning point, FLORES’s CoT was approximately 70% of that of B2W, and this highlights the superior efficiency of FLORES in navigating through the cluttered environment.
@@
TABLE II: Comparison of CoT for turning maneuvers
@@ V CONCLUSIONS
This work presents a novel wheel-legged robot, FLORES, featuring an innovative front-leg design that synthesizes the speed and efficiency of wheels with the versatility and adaptability of legs, which significantly enhances the performance on unstructured terrain. FLORES offers enhanced steering ability, and it is well-suited for dynamic hybrid locomotion in diverse scenarios with its high output-speed joint modules. Additionally, we propose an RL-based control framework, with a customized reward mechanism tailored to the unique mechanical design of FLORES. Through real-world experiments, FLORES demonstrates noteworthy flexibility and efficiency across various environments, validating the effectiveness of the proposed design and control strategy.